OpenStudy (anonymous):

solve 5x+2y=4 3x+4y+2z=6 7x+3y+4z=29

OpenStudy (anonymous):

I think you start with setting each of the equations equal to 29: 5x+2y=29 solve this one for y (in terms of x) 43x+4y+2z=29 67x+3y+4z=29 combine these two so that you eliminate z y= (29-5x)/2 and (86 - 67)x +(8-3)y = 19x+5y = 29. Now you could use these two equations to solve for x and y...but it's faster if you happen to have a graphing calculator. Graph the two lines and find their point of intersection.....(-4.94486, 26.86214) Now just plug these values in and solve for z....I leave that to you

OpenStudy (anonymous):

5x +2y =29 ------------(1) 43x + 4y + 2z = 29 -----(2) 67x + 3y + 4z = 29 -----(3) 2 times equation (2) gives 86x + 8y + 4z = 58 ------(4) equation (4) - equation (3) gives 19x + 5y = 29 --------(5) 5/2 times equation (1) gives 25/2 x + 5y = 29x5/2 ------(6) equation (5) - equation (6) gives (19-12.5)x = (29-5x29/2) therefore x = -6.69 substitute x value to equation (1) gives y = 31.225 and substitute x value and y value to equation (2) gives z = 95.885

OpenStudy (anonymous):

You would need to use the method of elimination 3x+4y+2z=6 7x+3y+4z=29 So then you could multiple the top by 3 and the bottom equation by -4 So then it would be 9x+12y+6z = 18 -28x-12y-16z=-116 then y would be out. and you would have -19x-19z = -98 . So then can you do it from there?

OpenStudy (anonymous):

then would i just plug in x and z into one of the equations to get y?

OpenStudy (anonymous):

Well you would need to manipulate from there, and find each value. You could solve for x and use it, or visa versa,

OpenStudy (anonymous):

okay thank you rick

OpenStudy (texaschic101):

7x + 3y + 4z = 29 3x + 4y + 2z = 6 -->(-2) --------------- 7x + 3y + 4z = 29 -6x - 8y - 4z = -12 (result of multiplying by -2) ---------------add x - 5y = 17 5x + 2y = 4 x - 5y = 17 -->(-5) ---------------- 5x + 2y = 4 -5x + 25y = - 85 (result of multiplying by -5) ---------------- 27y = - 81 y = -3 x - 5y = 17 x - 5(-3) = 17 x + 15 = 17 x = 17 - 15 x = 2 7x + 3y + 4z = 29 7(2) + 3(-3) + 4z = 29 14 - 9 + 4z = 29 5 + 4z = 29 4z = 29 - 5 4z = 24 z = 6 check... 3x + 4y + 2z = 6 3(2) + 4(-3) + 2(6) = 6 6 - 12 + 12 = 6 6 = 6 (correct) x = 2, y = -3 and z = 6